Suppose the nationwide high school dropout rate for 2014 was around 10.45%. If you checked the records of 462 students in high school in 2014, what is the probability that between 8.45% and 9.98% of them ended up dropping out?

Question 2 options:

1)

36.9864

2)

16.1006

3)

0.4506

4)

0.2906

5)

0.7094
Answer #1

Solution

Given that,

p = 0.1045

1 - p = 1 - 0.1045=0.8955

n = 462

\mu\hat p = p =0.1045

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] = \sqrt [(0.1045*0.8955) / 462] = 0.0142

= P( 0.0845 <\hat p<0.0998 )= P[(0.0845 - 0.1045) / 0.0142< ( \hat p - \mu \hat p ) / \sigma \hat p < (0.0998 - 0.1045) /0.0142 ]

= P(-1.4085< z <-0.3310)

= P(z <-0.33 ) - P(z < -1.41)

Using z table,   

= 0.3701 - 0.0795

=0.2906